Sally has two bags of peanuts; one with 74% peanuts, and the other with 27% peanuts. To make a 5lb bag of 36% peanuts, how much of each should Sally use?
Answer by addingup(3677)

0.74x+0.27(5-x) = 0.36(5)
0.74x+1.35-0.27x = 1.8
0.47x = 0.45
x = 0.95
You need 0.95 pounds of 74% and 5-0.95 = 4.05lbs of 0.27%